# D-convolution categories and Hopf algebras

### v1

@@ 1 Introduction
the category of D-modules on $G$, on $G/G_{ad}$ (the adjoint quotient of $G$ by itself) and the category of Harish-Chandra bimodules, respectively. These categories are interesting and important in geometric representation theory. Although the monoidal structures can be elegantly defined geometrically, their computations are usually difficult. Moreover, although it is known that the category ${\mathcal{D}}(G/G_{ad})$ has the structure of a braided tensor category [BD14], and is equivalent to the center of the other two monoidal categories [BITV23], the braiding is still difficult to compute in general.
@@
We then turn to braided monoidal structure of ${\mathcal{D}}_{\hbar}(G/G_{ad})^{\mathbb{C}^{\times}}$. This was first constructed in [BD14] using functoriality of $D$-modules. In the $\infty$-category setting, this was constructed in the work of [Ber17]. However, neither of these works gave an explicit description of the braiding. We apply Theorem 1.1 and show that a braiding exists on $A_{G/G_{ad}}$ in terms of the braiding of a double.
@@
The topological quantum double of $H_{G}$ and ${\mathcal{A}}_{G}$ are twist equivalent, and the double $D({\mathcal{A}}_{G})$ is a central extension of $A_{G/G_{ad}}$ by a primitive element $\delta^{*}$ in degree $-1$. The $R$-matrix of $D({\mathcal{A}}_{G})$ descends to an $R$-matrix in the quotient $A_{G/G_{ad}}=D({\mathcal{A}}_{G})/(\delta_{*})$.
@@
A twist $\theta$ exists for $A_{G/G_{ad}}$ (Remark 4.8 and Proposition A.11). In particular, the category of finite-dimensional modules of $A_{G/G_{ad}}$ has the structure of a pivotal braided monoidal category.
This theorem is proven in Section 4, Theorem 4.1. The fact that the double of $H_{G}$ and ${\mathcal{A}}_{G}$ are twist equivalent is expected from the work of [Ber17, BZG17]. It follows from the calculation of the $R$-matrix that when restricted to the heart of the standard t structure of D-module categories, the briading of Theorem 1.3 recovers the braiding of [BFO12] (see Section 4.1.1). We expect that the full braiding can be comparable to the one defined in [BD14]. We now give some explanation for the construction of the Hopf algebras in terms of topological quantum field theories.
@@ 1.1 Hopf algebras from 3 dimensional TQFTs
where the LHS is the derived category of ×$\mathbb{C}^{\times}\!\ltimes G({\mathcal{O}})$-equivariant $D$-modules on $\mathrm{Gr}_{G}$. Here $\mathbb{C}^{\times}$ acts by loop rotation. Turning-off the $\mathbb{C}^{\times}$ action on the LHS corresponds to $\hbar\to 0$ on the RHS. Taking the heart of the perverse t-structure on the LHS, one obtains the more familiar statement
@@
This TQFT is somewhat understood from geometric representation theory perspective. The category ${\mathcal{C}}$ is expected to be the category of D-modules on the character stack $G/G_{ad}$, which, using the work of [BZFN10, BZN12, BZGN17], can be written as
@@
Fortunately, we can completely control the braided tensor structure of ${\mathcal{C}}$ (at least on physical grounds) using quasi-triangular Hopf algebras. This was done in [DN24]. The idea is that the TQFT in question has a pair of transverse and complete boundary conditions (different from the cigar boundary condition). They are simply called the Dirichlet and Neumann boundary conditions. When this happens, it was argued in op.cit., that sandwiching the gauge theory by the two boundary conditions gives a fiber functor for the monoidal category ${\mathcal{C}}$ as well as the monoidal categories associated to the boundaries. For the TQFT in question, the Hopf algebra ${\mathcal{A}}_{G}$ and its double $D({\mathcal{A}}_{G})$ are precisely the symmetry algebras of the fiber functors, as in [DN24, Section 9]. Furthermore, the equivalence in Theorem 1.1 gives a consistency check of the proposal of [DN24] with the geometric approach of [BZFN10, BZN12]. [^1] We comment that this quotient by $\delta^{*}$ is a mild modification, and comes from our choice of treating the differential as an element in the algebra, similar to [BZN12].
@@
Unfortunately, besides the work of [Pim15], there hasn’t been much study of Lie bialgebras in cohomologically shifted settings. In [NP25], we studied 1-shifted Lie bialgebras and their quantizations. A 1-shifted Lie bialgebra is a graded Lie algebra $\mathfrak{h}$ with a coproduct
$\delta:\mathfrak{h}\to\mathrm{Sym}^{2}(\mathfrak{h})[1]$ (1.8)
satisfying a graded cocycle condition similar to ordinary Lie bialgebras. Such an object admits a graded 1-shifted dual $\mathfrak{h}^{*}[-1]$ and graded double $D(\mathfrak{h})=\mathfrak{h}\oplus\mathfrak{h}^{*}[-1]$. The coproduct $\delta_{D(\mathfrak{h})}:=(-\delta_{\mathfrak{h}},\delta_{\mathfrak{h}^{*}[-1]})$ on the double is a coboundary, using a 1-shifted version of $r$-matrix
$\delta_{D(\mathfrak{h})}=[\mathbf{r},\Delta],\qquad\mathbf{r}=\sum x_{i}\otimes c^{i}\in\mathfrak{h}\otimes\mathfrak{h}^{*}[-1].$ (1.9)
Here $\{x_{i}\}$ is a set of basis for $\mathfrak{h}$ and $c^{i}$ 1-shifted dual basis in $\mathfrak{h}^{*}[-1]$.
@@ 1.3 Organizations and conventions
In this paper, by “grading” it will always mean cohomological grading, whereas grading coming from any external $\mathbb{C}^{\times}$ action will simply be refered to as “weights”. Even though the algebras involved will be generated in various cohomological degrees (namely non-classical), one can perform a shearing using the $\mathbb{C}^{\times}$ grading so that they become classical. This shearing introduces extra signs in the coproduct formulas according to the Koszul sign rule, and we prefer not to deal with the signs.
@@ 2.1.1 Generalities on D-modules
[…]
@@ 2.1.2 Hecke categories
[…]
@@ 2.2.1 D-modules and de-Rham complexes
Let $X$ be a smooth affine scheme, and let $\Omega_{X}$ be the ring of differential operators on $X$. This graded algebra can be naturally identified with functions on $T[1]X$, the $(-1)$-shifted tangent space of $X$. This algebra admits a square-zero differential $d_{dR}$, the de-Rham differential. Let $\Omega_{X}^{dR}$ be the DG algebra given by $(\Omega_{X},d_{dR})$, which is the algebraic version of the de-Rham complex of $X$. There is a pair of adjoint functors between the DG categories of chain complexes
@@
where $\Omega$ sends a D-module $M$ to the de-Rham complex corresponding to $M$. One can show that the derived endomorphism algebra $\mathrm{REnd}({\mathcal{O}}_{X})$ is quasi-isomorphic to $\Omega_{X}^{dR}$, and the above adjunction is simply a tensor-Hom adjunction (hence Koszul duality). However, descending to the derived categories, this is usually not an equivalence. In [BD96], this was fixed by inverting morphisms of $\Omega_{X}^{dR}$ modules that are quasi-isomorphisms after applying $D$ (therefore not the usual derived category). Another way to deal with this is by passing to a $\mathbb{C}^{\times}$ equivariant category, and consider bounded derived categories, as in [BGS96]. We combine both approaches to obtain a Hopf-algebraic description of the categories.
For $X$ a smooth affine scheme, the algebra $D_{X}$ is a filtered algebra, whose associated graded is naturally identified with functions on $T^{*}X$. We consider the Rees algebra $D_{X}^{\hbar}$ associated to $D_{X}$, which is an algebra with a $\mathbb{C}^{\times}$ action, and the category of $\mathbb{C}^{\times}$-equivariant modules ${\mathcal{M}}_{\hbar}(X)^{\mathbb{C}^{\times}}$. Locally, the action of $\mathbb{C}^{\times}$ has degree $-1$ on the tangent vectors and $\hbar$. The formalism of the previous section works equally well in this setting, allowing us to define the derived category ${\mathcal{D}}_{\hbar}({\mathcal{X}})^{\mathbb{C}^{\times}}$ for a smooth stack ${\mathcal{X}}$, whose heart is the abelian category ${\mathcal{M}}_{\hbar}({\mathcal{X}})^{\mathbb{C}^{\times}}$.
On the other hand, the algebra $\Omega_{X}$ carries a natural $\mathbb{C}^{\times}$ action, whose weights coincide with the grading. Denote by ${\mathcal{A}}_{X}$ (following the notation of [BZN12]) the algebra
@@
[…]
This pair is induced by the $\mathbb{C}^{\times}$-equivariant version of the de-Rham complex. We call a morphism in $C({\mathcal{A}}_{X}\mathrm{-Mod}^{\mathbb{C}^{\times}})$ a $D$-quasiisomorphism if its image under $D_{\hbar}$ is a quasi-isomorphism. Denote by $D_{qs}({\mathcal{A}}_{X}\mathrm{-Mod}^{\mathbb{C}^{\times}})$ the localization of $K({\mathcal{A}}_{X}\mathrm{-Mod}^{\mathbb{C}^{\times}})$ (the homotopy category) with respect to $D$-quasiisomorphisms.
@@
Here $\xi_{x_{i}}^{R}(f)$ is the action of $\mathfrak{g}$ on ${\mathcal{O}}_{G}$ by left-invariant vector fields. Under the equivalence of equation (2.11), the monoidal structure of ${\mathcal{D}}_{\hbar}(G)^{\mathbb{C}^{\times}}$ corresponds to the group structure of $T[1]G=G\ltimes\mathfrak{g}[1]$, the extension of $G$ by the representation $\mathfrak{g}$ in degree $-1$. In the next section, we will write the corresponding coproduct more explicitly.
@@
Here $\xi_{b}^{ad}$ is the vector field on $G$ corresponding to $b$ (under the conjugation action), and $\langle$-,$\rangle$ - , - ⟩ is the natural pairing between vector fields and 1-forms. The term $\mathrm{Lie}_{b}$ is the Lie algebra action of $\mathfrak{g}$ on $M^{*}$. In the next sections, we will construct the coproduct corresponding to the monoidal structure under the equivalence of equation (2.11), as well as a braiding.
@@ 2.2.2 Harish-Chandra modules and Chevalley-Eilenberg complex
We now consider a similar Koszul duality statement for $\mathrm{HC}(\mathfrak{g},K)$, and again in the equivariant context. Let $U_{\hbar}(\mathfrak{g})$ be the Rees algebra associated to $U(\mathfrak{g})$. It is a $\mathbb{C}^{\times}$-equivariant algebra where $\mathbb{C}^{\times}$ acts on $\mathfrak{g}$ and $\hbar$ with weight $-1$. A vector space $M$ is called an asymptotic Harish-Chandra module, if it is a module of $U_{\hbar}(\mathfrak{g})$ and admits an action of $K$ such that $\hbar d\mathrm{act}(x)=\mathrm{Lie}_{x}$, where $d\mathrm{act}$ is the action of $\mathfrak{k}$ induced by the derivative of the $K$ action, and $\mathrm{Lie}_{x}$ is its action through the embedding $\mathfrak{k}\to\mathfrak{g}$. Since the relation $\hbar d\mathrm{act}(x)=\mathrm{Lie}_{x}$ is equivariant with respect to the $\mathbb{C}^{\times}$ action, one can define the category $\mathrm{HC}_{\hbar}(\mathfrak{g},K)^{\mathbb{C}^{\times}}$ to be the category of equivariant asymptotic Harish-Chandra modules.
@@
[…]
Call a morphism of ${\mathcal{A}}_{\mathfrak{g},K}$ modules a $\mathfrak{g}$-quasiisomorphism if its image under $D_{\mathfrak{g},K}$ is a quasi-isomorphism, and denote by $D_{qs}({\mathcal{A}}_{\mathfrak{g},K}\mathrm{-Mod}^{\mathbb{C}^{\times}})$ the localization of $K({\mathcal{A}}_{\mathfrak{g},K}\mathrm{-Mod}^{\mathbb{C}^{\times}})$ with respect to $\mathfrak{g}$-quasiisomorphisms.
@@
[…]

### v2

@@ 1 Introduction
the category of D-modules on $G$, on $G/G_{ad}$ (the adjoint quotient of $G$ by itself) and the category of Harish-Chandra bimodules, respectively. These categories are interesting and important in geometric representation theory. Although the monoidal structures can be elegantly defined geometrically, their computations are usually difficult. Moreover, although it is known that the category ${\mathcal{D}}(G/G_{ad})$ has the structure of a braided tensor category [BD14], and is equivalent to some notion of the center of the other two monoidal categories [BIT+23], the braiding is still difficult to compute in general.
@@
We then turn to braided monoidal structure of ${\mathcal{D}}_{\hbar}(G/G_{ad})^{\mathbb{C}^{\times}}$. This was first constructed in [BD14, Appendix B] using functoriality of $D$-modules [^1] Although in this work the authors worked with the category of constructible sheaves, their construction works equally well for the category $\mathcal{D}$$_{\hbar}$, as all the functors needed for the construction have counterparts in the asymptotic setting.. In the $\infty$-category setting, this was constructed in the work of [BER17]. However, neither of these works gave an explicit description of the braiding. We apply Theorem 1.1 and show that a braiding exists on $A_{G/G_{ad}}$ in terms of the braiding of a double.
@@
The topological quantum double of $H_{G}$ and ${\mathcal{A}}_{G}$ are twist equivalent, and the double $D({\mathcal{A}}_{G})$ is a central extension of $A_{G/G_{ad}}$ by a primitive element $\delta^{*}$ in degree 1-1. The $R$-matrix of $D({\mathcal{A}}_{G})$ descends to an $R$-matrix in the quotient $A_{G/G_{ad}}=D({\mathcal{A}}_{G})/(\delta_{*})$.
@@
A ribbon element $\theta$ exists for $A_{G/G_{ad}}$ (Remark 4.8 and Proposition A.11). In particular, the category of finite-dimensional modules of $A_{G/G_{ad}}$ has the structure of a pivotal braided monoidal category.
This theorem is proven in Section 4, Theorem 4.1. The fact that the double of $H_{G}$ and ${\mathcal{A}}_{G}$ are twist equivalent is expected from the work of [BER17, BG17]. It follows from the calculation of the $R$-matrix that when restricted to the heart of the standard t structure of D-module categories, the braiding of Theorem 1.3 recovers the braiding of [BFO12] (see Section 4.1.1). We expect that the full braiding can be comparable to the one defined in [BD14]. We now give some explanation for the construction of the Hopf algebras in terms of topological quantum field theories.
@@ 1.1 Hopf algebras from 3 dimensional TQFTs
where the LHS is the derived category of $\mathbb{C}^{\times}\!\ltimes G({\mathcal{O}})$-equivariant $D$-modules on $\mathrm{Gr}_{G}$. Here $\mathbb{C}^{\times}$ acts by loop rotation. Turning-off the $\mathbb{C}^{\times}$ action on the LHS corresponds to $\hbar\to 0$ on the RHS. Taking the heart of the perverse t-structure on the LHS, one obtains the more familiar statement
@@
This TQFT is somewhat understood from geometric representation theory perspective. The category $\mathcal{C}$} is expected to be the category of D-modules on the character stack $G/G_{ad}$, which, using the work of [BFN10, BN12, BGN17], can be written as
@@
Fortunately, we can completely control the braided tensor structure of $\mathcal{C}$} (at least on physical grounds) using quasi-triangular Hopf algebras. This was done in [DN24]. The idea is that the TQFT in question has a pair of transverse and complete boundary conditions (different from the cigar boundary condition). They are simply called the Dirichlet and Neumann boundary conditions. When this happens, it was argued in op.cit., that sandwiching the gauge theory by the two boundary conditions gives a fiber functor for the monoidal category $\mathcal{C}$} as well as the monoidal categories associated to the boundaries. For the TQFT in question, the Hopf algebra ${\mathcal{A}}_{G}$ and its double $D({\mathcal{A}}_{G})$ are precisely the symmetry algebras of the fiber functors, as in [DN24, Section 9]. Furthermore, the equivalence in Theorem 1.1 gives a consistency check of the proposal of [DN24] with the geometric approach of [BFN10, BN12]. [^2] We comment that this quotient by $\delta^{*}$ is a mild modification, and comes from our choice of treating the differential as an element in the algebra, similar to [BN12].
@@
Unfortunately, besides the work of [PIM15], there hasn’t been much study of Lie bialgebras in cohomologically shifted settings. In [NP25], we studied 1-shifted Lie bialgebras and their quantizations. A 1-shifted Lie bialgebra is a graded Lie algebra $\mathfrak{a}$ with a cobracket $\delta$ of degree $1$, satisfying a graded cocycle condition similar to ordinary Lie bialgebras. It was show in loc.cit that such an object admits a 1-shifted dual $\mathfrak{a}^{*}[-1]$ and double $D(\mathfrak{a})$, and that the cobracket on the double is coboundary. The double $D(\mathfrak{a})$ is a 1-shifted metric Lie algebra. We will give a quick review of these in Section 3.1.
@@ 1.3 Organizations and conventions
In this paper, by “grading” it will always mean cohomological grading, whereas grading coming from any external $\mathbb{C}^{\times}$ action will simply be refered to as “weights”. Even though the algebras involved will be generated in various cohomological degrees (namely non-classical), one can perform a shearing using the $\mathbb{C}^{\times}$ grading so that they become classical. This shearing introduces extra signs in the coproduct formulas according to the Koszul sign rule, and we prefer not to deal with the signs. In particular, Koszul sign rule is always assumed, and all Hopf algebras are Hopf algebras in the category of chain complexes.
@@ 2.1.1 Generalities on D-modules
${\mathcal{D}}(K\setminus G)^{K}$$D(\mathrm{HC}(\mathfrak{g},K)).$$\scriptstyle{\Gamma}$ (2.3)
@@ 2.1.2 Hecke categories
We can consider 2 specializations. The first one is when $K=1$ the trivial group, in which case the Hecke category is simply ${\mathcal{D}}(G)$. We can also consider the case when $G=K\times K$ and $K=K_{\Delta}$ the diagonal group, then the stack $K\setminus G/K$ is naturally identified with $K/K_{ad}$, the adjoint quotient of $K$ by $K$. This stack is called the character stack of $K$, and can be identified with the topological loop space of /K*/K, as in [BN12]. From the above discussions, the category ${\mathcal{D}}(K/K_{ad})$ admits a monoidal structure, via the identification with $K\setminus G/K$. Moreover, it is braided monoidal [BD14].
@@ 2.2.1 D-modules and de-Rham complexes
Let $X$ be a smooth affine scheme, and let $\Omega_{X}$ be the ring of differential forms on $X$. This graded algebra can be naturally identified with functions on $T[1]X$, the (1)(-1)-shifted tangent space of $X$. This algebra admits a square-zero differential $d_{dR}$, the de-Rham differential. Let $\Omega_{X}^{dR}$ be the DG algebra given by $(\Omega_{X},d_{dR})$, which is the algebraic version of the de-Rham complex of $X$. There is a pair of adjoint functors between the DG categories of chain complexes
@@
where $\Omega$ sends a D-module $M$ to the de-Rham complex corresponding to $M$. One can show that the derived endomorphism algebra $\mathrm{REnd}({\mathcal{O}}_{X})$ is quasi-isomorphic to $\Omega_{X}^{dR}$, and the above adjunction is simply a tensor-Hom adjunction (hence Koszul duality). However, descending to the derived categories, this is usually not an equivalence. In [BD96, Section 7.2.5], this was fixed by inverting morphisms of $\Omega_{X}^{dR}$ modules that are quasi-isomorphisms after applying $D$ (therefore not the usual derived category). Another way to deal with this is by passing to a $\mathbb{C}^{\times}$ equivariant category, and considering bounded derived categories, as in [BGS96]. This has the advantage that the Koszul dual is a ordinary Hopf algebra rather than a DG Hopf algebra (essentially treating the de-Rham differential as a generator as in [BN12, Section 5]), but the drawback in our setting is that the bounded derived category does not admit a monoidal structure, as the map $G\times G\to G$ is not proper. We therefore combine the two approaches: we consider $\mathbb{C}^{\times}$ equivariant categories that are not neccesarily bounded, and invert quasi-isomorphisms in the same way as [BD96, Section 7.2.5].
For $X$ a smooth affine scheme, the algebra $D_{X}$ is a filtered algebra, whose associated graded is naturally identified with functions on $T^{*}X$. We consider the Rees algebra $D_{X}^{\hbar}$ associated to $D_{X}$, which is an algebra with a $\mathbb{C}^{\times}$ action, and the category of $\mathbb{C}^{\times}$-equivariant modules ${\mathcal{M}}_{\hbar}(X)^{\mathbb{C}^{\times}}$. Locally, the action of $\mathbb{C}^{\times}$ has degree 1-1 on the tangent vectors and $\hbar$. The formalism of the previous section works equally well in this setting, allowing us to define the derived category ${\mathcal{D}}_{\hbar}({\mathcal{X}})^{\mathbb{C}^{\times}}$ for a smooth stack $\mathcal{X}$}, whose heart is the abelian category ${\mathcal{M}}_{\hbar}({\mathcal{X}})^{\mathbb{C}^{\times}}$.
On the other hand, the algebra $\Omega_{X}$ carries a natural $\mathbb{C}^{\times}$ action, whose weights coincide with the grading. Denote by ${\mathcal{A}}_{X}$ (following the notation of [BN12]) the algebra
@@
$C({\mathcal{M}}_{\hbar}(X)^{\mathbb{C}^{\times}})$$C({\mathcal{A}}_{X}\mathrm{-Mod}^{\mathbb{C}^{\times}})$$\scriptstyle{\Omega_{\hbar}}$$\scriptstyle{D_{\hbar}}$ (2.9)
This pair is induced by the $\mathbb{C}^{\times}$-equivariant version of the de-Rham complex. More explicitly, consider the object
$DR_{\hbar,X}:=D_{X}^{\hbar}\otimes_{{\mathcal{O}}_{X}}{\mathcal{A}}_{X},\qquad d=\sum_{i}(v_{i})_{L}\otimes(\omega^{i})_{L}+(\hbar)_{L}\otimes(\delta_{dR})_{L},$ (2.10)
where $\{v_{i}\}$ is a (local) basis for $T_{X}$ and $\{\omega^{i}\}$ a dual basis for $\Omega^{1}(X)$, and the subscrip $L$ means left multiplications. Note the tensor is over ${\mathcal{O}}_{X}$ as left modules. This is a $D_{X}^{\hbar,op}-{\mathcal{A}}_{X}$ bimodule, and the functors in equation (2.9) are given by
$\Omega_{\hbar}(M):=\mathrm{Hom}_{D_{X}^{\hbar}}(DR_{\hbar,X},M),\qquad D_{\hbar}(N):=DR_{\hbar,X}\otimes_{{\mathcal{A}}_{X}}N.$ (2.11)
We call a morphism in $C({\mathcal{A}}_{X}\mathrm{-Mod}^{\mathbb{C}^{\times}})$ a $D$-quasiisomorphism if its image under $D_{\hbar}$ is a quasi-isomorphism. Denote by $D_{qs}({\mathcal{A}}_{X}\mathrm{-Mod}^{\mathbb{C}^{\times}})$ the localization of $K({\mathcal{A}}_{X}\mathrm{-Mod}^{\mathbb{C}^{\times}})$ (the homotopy category) with respect to $D$-quasiisomorphisms. By repeating the proof of [BD96, Lemma 7.2.4 & Section 7.6.11] in the $\mathbb{C}^{\times}$-equivariant setting, one arrives at the following result.
@@
Proof.
This is a repeat of the proof of [BD96, Lemma 7.2.4]. Indeed, one shows that for any $M\in C({\mathcal{M}}_{\hbar}(X)^{\mathbb{C}^{\times}})$, the canonical adjunction $D\Omega M\to M$ is a quasi-isomorphism. Therefore, the adjunction $N\to\Omega DN$ is a $D$-quasiisomorphism. The second part of the statement follows from [BD96, Section 7.6.11], again applied in the $\mathbb{C}^{\times}$-equivariant setting.
@@
Here $\xi_{x_{i}}^{R}(f)$ is the action of $\mathfrak{g}$ on ${\mathcal{O}}_{G}$ by left-invariant vector fields. Under the equivalence of equation (2.13), the monoidal structure of ${\mathcal{D}}_{\hbar}(G)^{\mathbb{C}^{\times}}$ corresponds to the group structure of $T[1]G=G\ltimes\mathfrak{g}[1]$, the extension of $G$ by the representation $\mathfrak{g}$ in degree 1-1. In the next section, we will write the corresponding coproduct more explicitly.
@@
Here $\xi_{b}^{ad}$ is the vector field on $G$ corresponding to $b$ (under the conjugation action), and $\langle-,$-$\rangle$ is the natural pairing between vector fields and 1-forms. The term $\mathrm{Lie}_{b}$ is the Lie algebra action of $\mathfrak{g}$ on $M^{*}$. In the next sections, we will construct the coproduct corresponding to the monoidal structure under the equivalence of equation (2.13), as well as a braiding.
@@ 2.2.2 Harish-Chandra modules and Chevalley-Eilenberg complex
We now consider a similar Koszul duality statement for $\mathrm{HC}(\mathfrak{g},K)$, and again in the equivariant context. Let $U_{\hbar}(\mathfrak{g})$ be the Rees algebra associated to $U(\mathfrak{g})$. It is a $\mathbb{C}^{\times}$-equivariant algebra where $\mathbb{C}^{\times}$ acts on $\mathfrak{g}$ and $\hbar$ with weight 1-1. A vector space $M$ is called an asymptotic Harish-Chandra module, if it is a module of $U_{\hbar}(\mathfrak{g})$ and admits an action of $K$ such that $\hbar d\mathrm{act}(x)=\mathrm{Lie}_{x}$, where $d\mathrm{act}$ is the action of $\mathfrak{k}$ induced by the derivative of the $K$ action, and $\mathrm{Lie}_{x}$ is its action through the embedding $\mathfrak{k}\to\mathfrak{g}$. Since the relation $\hbar d\mathrm{act}(x)=\mathrm{Lie}_{x}$ is equivariant with respect to the $\mathbb{C}^{\times}$ action, one can define the category $\mathrm{HC}_{\hbar}(\mathfrak{g},K)^{\mathbb{C}^{\times}}$ to be the category of $\mathbb{C}^{\times}$-equivariant asymptotic Harish-Chandra modules.
@@
$C(\mathrm{HC}_{\hbar}(\mathfrak{g},K)^{\mathbb{C}^{\times}})$$C({\mathcal{A}}_{\mathfrak{g},K}\mathrm{-Mod}^{\mathbb{C}^{\times}})$$\scriptstyle{\Omega_{\mathfrak{g},K}}$$\scriptstyle{D_{\mathfrak{g},K}}$ (2.20)
Call a morphism of ${\mathcal{A}}_{\mathfrak{g},K}$ modules a $\mathfrak{g}$-quasiisomorphism if its image under $D_{\mathfrak{g},K}$ is a quasi-isomorphism, and denote by $D_{qs}({\mathcal{A}}_{\mathfrak{g},K}\mathrm{-Mod}^{\mathbb{C}^{\times}})$ the localization of $K({\mathcal{A}}_{\mathfrak{g},K}\mathrm{-Mod}^{\mathbb{C}^{\times}})$ with respect to $\mathfrak{g}$-quasiisomorphisms. The following, as Proposition 2.4, is a repeat of [BD96, Section 7.7.7] in the $\mathbb{C}^{\times}$-equivariant setting.
@@
$C(\mathrm{HC}_{\hbar}(\mathfrak{g},K)^{\mathbb{C}^{\times}})$$C({\mathcal{A}}_{\mathfrak{h},K}\mathrm{-Mod}^{\mathbb{C}^{\times}})$$\scriptstyle{\Omega_{\mathfrak{h},K}}$$\scriptstyle{D_{\mathfrak{h},K}}$ (2.22)